OpenStudy (fellowroot):

Find delta >0 The rest of problem in the pic

OpenStudy (anonymous):

I do not see a picture.

OpenStudy (fellowroot):

I am providing the question as well as the books solution. My question is just help me understand whats going on in the solution. where do they get 4 from???

OpenStudy (anonymous):

deltax?

OpenStudy (fellowroot):

pic included

OpenStudy (perl):

from the line above

OpenStudy (perl):

| x- 5 | = | x - 2 - 3 | <= |x -2 | + |-3| by triangle inequality theorem

OpenStudy (fellowroot):

okay i got the triangle inequality, but dont get the 4 or the 1/12

OpenStudy (perl):

| x- 5 | = | x - 2 - 3 | <= |x -2 | + |-3| by triangle inequality theorem < 1 / 12 + 3 = 3.08333.. < 4

OpenStudy (fellowroot):

could i have picked different numbers other than 1/12 and 4

OpenStudy (perl):

yes, but you would end up with a number less than 1/3 , in the next step

OpenStudy (fellowroot):

is that bad

OpenStudy (perl):

which is fine

OpenStudy (fellowroot):

what about instead of saying 4 say 10 or 100 or 1000. cant i just pick some really big number to be safe

OpenStudy (perl):

no you dont want bigger than 4, because the next step has to be less than 1/3

OpenStudy (fellowroot):

does it have to be less than 1/3 or exactly = to 1/3

OpenStudy (perl):

What happens if we use a less tight bound if you had | x - 5 | < 5 , then the next step would be | x^2 -7x + 10| = | x - 5| | x - 2 | < 1/12 * 5 = 5/12, which is greater than 1/3

OpenStudy (fellowroot):

so you cant pick 5. got ya

OpenStudy (perl):

the biggest you can choose is 4, the least value is 3 + 1/12

OpenStudy (fellowroot):

thank you. i will now try some of the other problems like this. this really helped!